### v1

# Uni-NaVid: A Video-based Vision-Language-Action Model for Unifying Embodied Navigation Tasks

Abstract
A practical navigation agent must be capable of handling a wide range of interaction demands, such as following instructions, searching objects, answering questions, tracking people, and more. Existing models for embodied navigation fall short of serving as practical generalists in the real world, as they are often constrained by specific task configurations or pre-defined maps with discretized waypoints. In this work, we present Uni-NaVid, the first video-based vision-language-action (VLA) model designed to unify diverse embodied navigation tasks and enable seamless navigation for mixed long-horizon tasks in unseen real-world environments. Uni-NaVid achieves this by harmonizing the input and output data configurations for all commonly used embodied navigation tasks and thereby integrating all tasks in one model. For training Uni-NaVid, we collect 3.6 million navigation data samples in total from four essential navigation sub-tasks and foster synergy in learning across them. Extensive experiments on comprehensive navigation benchmarks clearly demonstrate the advantages of unification modeling in Uni-NaVid and show it achieves state-of-the-art performance. Additionally, real-world experiments confirm the model’s effectiveness and efficiency, shedding light on its strong generalizability.
Figure 1: Uni-NaVid learns general navigation skills across four embodied navigation tasks using 3.6 million navigation samples. Uni-NaVid only takes online RGB video frames and language instructions as input and output actions, achieving general navigation ability in a real-world deployment.
00footnotetext: First author (e-mail: zhngjizh@gmail.com). $\dagger$ Corresponding authors (e-mail: zhangzz@galbot.com, hewang@pku.edu.cn).
1 Introduction
Embodied navigation is fundamental for supporting various interaction demands. Recently, instruction-following navigation techniques have advanced significantly with the development of multimodal models. However, existing studies [2, 79, 98, 51, 100, 68, 96, 66, 61, 94] either develop individual models for different tasks, e.g., vision-and-language navigation [37, 39], object goal navigation [11], embodied question answering [18], following [97, 30, 56], etc, or explore task unification in a simplified setting (Tab. 1). They all fall short as full-stack navigation models qualified for practical general-purpose deployment.
Methods Action Embodied Navigation Tasks
D.E. C.E. VLN [37] ObjNav [64] EQA [18] Follow [56]
VLMaps [28]
NaviLLM [98]
InstructNav [51]
Poliformer [90]
Uni-NaVid
Table 1: Task and setting comparison. Uni-NaVid is developed to address four embodied navigation tasks, generating action outputs in continuous environments. C.E.: Continuous Environment; D.E.: Discrete Environment.
Developing a full-stack navigation model poses significant challenges, not only in unifying task modeling but also in integrating heterogeneous data in joint use. Initial efforts adopt imitation learning (IL) [72, 78, 55] or reinforcement learning (RL) [90, 82] to learn general navigation skills in simulation environments or limited diverse real-world environments. However, due to the limited rendering quality and diversity of simulators, these approaches often encounter the “sim-to-real” gap and suffer from poor generalization across diverse navigation tasks [22, 5, 33]. Recent studies [100, 98, 51, 50, 67] have attempted to achieve a higher degree of unification on top of pre-trained LLMs, but they simplify the problem to some extent by adopting discretized modeling approaches. They rely on pre-defined graphs for learning decision-making, thus sacrificing the output flexibility and resulting in a low upper bound, especially for real-world deployment.
In this work, we build Uni-NaVid, a Vision-Language-Action (VLA) model for unifying diverse commonly demanded navigation tasks (Tab. 1). It takes video streams and natural language as input, adaptively generates actions following instructions, and automatically responds to questions if raised in an end-to-end manner. We make the first endeavor to address three essential questions, as elaborated below, in pursuit of ultimate real-world application.
How to unify different navigation tasks with minimal problem simplification and maximum utilization of data? We target an up-to-date unification modeling in Uni-NaVid where it takes single-view egocentric video streams and natural language as inputs, and end-to-end generate actions or linguistic responses for instructions or questions adaptively. We follow the natural characteristic of Large Language Models (LLMs) to cross-modally interleave the inputs and outputs of different navigation tasks into a single sequence in the token space to enable joint learning with next-token-prediction paradigm for Uni-NaVid. It incorporates four widely demanded navigation tasks, including Vision-and-Language Navigation (VLN), object-goal navigation, embodied question answering, and human following. The corresponding training data are all collected from simulation environments. Besides, we further incorporate Internet real-world data for Video Question Answering (VQA) [7, 44] and video captioning [16] as auxiliary tasks into this unified model to enhance navigation performance and promote sim-to-real generalization. This is a groundbreaking attempt in this field at unified cross-modal, cross-task, and cross-sim-to-real modeling. It achieves impressive synergy in a broad range and results in a full-stack navigation agent.
How to improve the efficiency of such a unification design to enable a deployment-friendly model? In addition to unification modeling, we also prioritize efficient model design for practical deployment. Specifically, we adopt an online token merging mechanism to compress near historical frames with a relatively low ratio while compressing far historical frames with a relatively high ratio. In this way, Uni-NaVid learn compact representations that maintain not only fine-grained spatial information but also structured temporal information, thus speeding up the model inference by reducing the token number. Besides, Uni-NaVid adopts a foresight prediction to generate actions for a future horizon at once instead of step-by-step. This allows us to use an asynchronous strategy for action inference and execution.
How to verify the synergy between different tasks? We conduct comprehensive ablation studies to verify the synergy over different navigation tasks, as well as between simulation and real-world tasks in both simulation and real-world environments. We find they promote each other with unification modeling in such a broad range and help Uni-NaVid attain clear superiority in all sub-tasks [37, 64, 18, 31] compared to recent works. Besides, we also evaluate Uni-NaVid on long-horizon real-world tasks, which include multiple subtasks, and witness it can complete them in a zero-shot manner as shown in Figure 1.
By exploring these three essential questions, we demonstrate the advantages of unified modeling and joint data utilization across different tasks and in sim-to-real transfer. We hope this will foster future progress in developing an up-to-date full-stack navigation agent for general-purpose real-world applications.
2 Related Works
Figure 2: Pipeline of Uni-NaVid. Our method takes only single-view RGB frames $\{\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\}$ and a natural language instruction $\mathcal{I}$ as input. For each frame, we extract 64 visual tokens using the vision encoder and then use online token merging to accelerate the model while retaining compact visual information. The merged tokens and instruction tokens are sent to the large language model to obtain actions for navigation or answers for embodied question-answering.
Generalist Embodied Navigation. Embodied navigation [2, 79, 96] requires agents to navigate in unseen environments based on human instructions. There is extensive literature on embodied navigation; here, we focus on four mainstream tasks that involve both visual information and language instructions: Vision-and-Language Navigation [4, 37, 39], Object Goal Navigation [11], Embodied Question Answering [18], and Human Following [31, 56]. Early efforts [72, 78, 55, 82] towards a generalist-embodied navigation model involved multi-task navigation datasets and directly learning navigation skills, showing initial success in multi-task performance. However, these methods experienced performance drops when deployed in novel environments, especially in real-world settings. In recent years, advanced approaches [98, 51, 29, 50, 101, 67] have leveraged the generalization capabilities of large language models to improve multi-task navigation. These models show promising generalizability across navigation tasks but rely on extensive prompting, which impacts time efficiency. In contrast, our video-based large language model is trained end-to-end for multi-task navigation, offering robust generalization and computational efficiency for tasks like human following.
Large Language Models for Navigation. Large Language Models (LLMs)[17, 46, 102] have been introduced into robotic navigation due to their generalization capabilities in understanding and planning. One straightforward approach[100, 51, 50, 67] is to use off-the-shelf large language models in a zero-shot manner. These methods employ visual foundation models [19, 46] to describe surrounding environments in text format, prompting the language model to select landmarks that guide the agent. However, abstracting dense visual information into text and relying on discrete landmarks results in sparse environmental observations and is limited to static environments. Another approach [95, 90] trains a video-based large language model end-to-end with low-level actions to enable continuous movement. However, it faces efficiency challenges in long-horizon tasks. In contrast, Uni-NaVid implements an online visual token merging strategy, optimizing training efficiency for long-horizon tasks and supporting non-blocking execution in real-world environments.
3 Method
[…]
Overview. As shown in Figure 2, the architecture of Uni-NaVid is designed for general-purpose navigation which is composed of two main components: a vision encoder and a large language model (LLM). First, the video is encoded into a sequence of visual tokens via a vision encoder. The visual tokens are spatially and temporally merged, resulting in a compact sequence of visual tokens. Then they are projected into a space aligned with language tokens, which are referred to as observation tokens. As common, the instructions are also tokenized as a set of tokens, called instruction tokens. Both the observation and instruction tokens are concatenated and passed to the LLM, which infers 4 action tokens representing the next four actions.
3.1 Model of Uni-NaVid
Observation encoding. Given the ego-centric video up to time $T$, denoted by $\mathcal{O}_{T}=\{\mathbf{x}_{1}.\cdots,\mathbf{x}_{T}\}$, we encode the video to a sequence of observation tokens. For each frame $\mathbf{x}_{t}$, we first get its visual tokens $\mathbf{X}_{t}\in\mathbb{R}^{N_{x}\times C}$ with a vision encoder (EVA-CLIP [70] in implementation), where $N_{x}$ is the patch number ($N_{x}$ is set to 256) and $C$ is the embedding dimension. Inspired by vision-language models [46, 44], we also perform a vision-language alignment by projecting the visual feature embedding into the language space, which denotes as observation tokens $\mathbf{E}^{V}_{t}\in\mathbb{R}^{N_{x}\times C}$.
$\mathbf{E}^{V}_{t}=P_{V}(\mathbf{X}_{t}),$ (1)
Where $P_{V}$ represents the cross-modality projector. Although observation tokens provide crucial visual information for the agent to plan its next actions, the linearly increasing number of these tokens $(T\times N_{x})$ leads to progressively longer processing times, making it impractical to use them directly for navigation purposes.
Online visual token merging. To balance the efficiency and effectiveness of visual tokens, we design a token merging strategy to retain navigation-relevant visual information for Uni-NaVid. This strategy builds on the key navigation insight that recent observations are more critical, and that visual information between consecutive frames and neighboring pixels could be redundant.
Drawing inspiration from the Atkinson-Shiffrin memory model [6, 69], we categorize visual tokens into current visual tokens $\mathbf{X}_{\text{curr}}$, short-term visual tokens $\mathbf{X}_{\text{short}}$, and long-term visual tokens $\mathbf{X}_{\text{long}}$. These tokens are grouped based on their timestamps relative to the most recent frame. Current visual tokens capture sufficient visual information for the agent to perceive its immediate surroundings and to plan subsequent paths. Short-term visual tokens maintain temporal continuity, preserving relevant information momentarily to support ongoing planning. Long-term visual tokens facilitate the integration and retrieval of knowledge across the entire navigation history. For each category of visual tokens, we apply a grid pooling operation at different pooling resolutions:
$\mathbf{X}_{1:T}=\begin{cases}\mathbf{X}_{\text{curr}}=GridPool(\mathbf{X}_{t},\alpha_{\text{curr}}),&\text{if t = T}\\ \mathbf{X}_{\text{short}}=GridPool(\mathbf{X}_{t},\alpha_{\text{short}}),&\text{if t $\in$ [T-B, T)}\\ \mathbf{X}_{\text{long}}=GridPool(\mathbf{X}_{t},\alpha_{\text{long}}),&\text{if t $\in$ [1, T-B)}\end{cases}$ (2)
where GridPool(·) is a grid pooling operation  [44, 95], squeezing the tokens from $N_{x}$ to $\frac{N_{x}}{\alpha^{2}}$, and $B$ (set to $64$) is the length of the buffer of shorter memory. Here, we adopt the $\alpha_{\text{curr}}=2$, $\alpha_{\text{short}}=8$, $\alpha_{\text{long}}=16$, leads to visual tokens as $\mathbf{X}_{\text{curr}}\in\mathbb{R}^{64\times C}$, $\mathbf{X}_{\text{short}}\in\mathbb{R}^{4\times C}$, $\mathbf{X}_{\text{long}}\in\mathbb{R}^{1\times C}$, respectively. These hyperparameters are determined through empirical experimentation and can be adjusted if memory and computational resources are not primary constraints.
In the navigation process, the agent consistently observes new frames. When a new frame at time $T+1$ arrives, we apply grid pooling to the most recent visual tokens at time $T$ and the oldest short-term visual tokens at time $T-B$, then insert them into the short-term and long-term visual tokens, respectively:
$\mathbf{X}_{\text{curr}\rightarrow\text{short}}$ $=GridPool(\mathbf{X}_{\text{curr}},\frac{\alpha_{\text{short}}}{\alpha_{\text{curr}}}),$ (3)
$\mathbf{X}_{\text{short}\rightarrow\text{long}}$ $=GridPool(\mathbf{X}_{\text{short}},\frac{\alpha_{\text{long}}}{\alpha_{\text{short}}}).$ (4)
In this way, we can online update the short-term visual tokens and long-term visual tokens with previously generated tokens. Furthermore, we compare the cosine similarity of the $\mathbf{X}_{\text{short}\rightarrow\text{long}}$ with the latest long-term visual tokens $\mathbf{X}_{\text{long}}$ at time $T-B-1$. If the similarity is above a given threshold $\tau$, we merge them based on the number of frames previously merged in the latest long-term visual tokens, $K$.
$\mathbf{X}_{\text{long}}$ $=\frac{1}{K+1}\left(K\mathbf{X}_{\text{long}}+\mathbf{X}_{\text{short}\to\text{long}}\right),$ (5)
subject to $\cos\left(\mathbf{X}_{\text{long}},\mathbf{X}_{\text{short}\to\text{long}}\right)>\tau.$ (6)
We insert new long-term visual tokens $\mathbf{X}_{\text{short}\to\text{long}}$ if their similarity falls below $\tau$ (empirically set to $\tau=0.95$ [69]), which indicates they contain relatively distinct visual information. This fusion in temporal preserves the observation history in a highly compact form (in a length of $M\ll T-B-1$), with minimal computational cost for merging these tokens. Only frames at the boundaries of the visual token types require computation through parallel pooling operations, making the process computationally efficient and naturally suited for online deployment in real-world navigation applications. Compared to existing video-based large language models [95, 69, 44], this merging strategy significantly reduces processing time, achieving an average of 0.2 seconds per inference. This improvement becomes increasingly notable when handling longer video sequences. A detailed analysis of time efficiency is provided in the Supplementary Materials.
Action planning. For action planning, we follow the existing work [95] and use special tokens <NAV> to inform the LLM to predict navigation actions. Accordingly, we concatenate the observation token sequence $\mathbf{E}^{V}_{1:T}$ (obtained through Eq. 1 with visual tokens of Eq. 2) with special token <NAV> and instruction tokens and input them into the LLM to infer 4 action tokens $\{\mathbf{E}^{A}_{T},\cdots,\mathbf{E}^{A}_{T+3}\}$:
Input: $\{Long\_term\_frames\}\{Shot\_term\_frames\}\\ \{Current\_frame\}<\!NAV\!>\{Instruction\}$ Output: $<\!Action\_0\!>\!<\!Action\_1\!>\!<\!Action\_2\!>\!\\ \!<\!Action\_3\!>$
[…]
3.2 Training of Uni-NaVid
Figure 3: Visualization of training data. We visualize the combination of training data (5.9M), video frame counts, and the most common words in navigation instructions.
To train a large-scale model for generalist navigation, it is essential to gather extensive and diverse navigation data across various tasks and environments. However, directly collecting vast amounts of real-world navigation data can be cost-prohibitive. To address this, we propose two key approaches for training Uni-NaVid: first, we collect multi-task navigation data in a wide range of synthetic environments (861 scenes in total), enabling Uni-NaVid to acquire general navigation skills; second, we co-tune Uni-NaVid with real-world video-based question-answering data, enhancing its understanding of real-world images and supporting its open-vocabulary knowledge.
Multi-task navigation data. We collect the largest multi-task navigation dataset to date on the Habitat simulator environment [64], including 3.6M samples across four different navigation tasks introduced below. They are all curated in a unified format. Details are in Supplementary.
(A) Vision-and-language navigation [37] requires the agent to comprehend language instructions in order to follow the described path. We collect two types of instructions: one type uses landmarks and describes required motions between them, while the other specifies low-level actions, such as moving forward by 5 steps. We collect 2.4M samples based on both VLN-CE R2R [37] and RxR [40].
(B) Object goal navigation [64]  involves the agent navigating an environment to locate a specific object based on provided visual or linguistic cues. This task assesses the agent’s ability to perceive objects, understand spatial relationships, and execute efficient search strategies. We gather 483k samples from datasets on Habitat Matterport 3D dataset [57].
(C) Embodied question answering [18] requires the agent to navigate to the related area for question answering. It involves spatial reasoning, object description, and understanding contextual information, requiring the ability to integrate perception, language comprehension, and decision-making. Following the setup in EQA [18], the agent first navigates to the target related to the question, issues a stop action, and then provides an answer. We collect 240k video-action samples and 10k video-answering samples on the EQA dataset [18] on Matterport 3D environments [10].
(D) Human following [31] requires the agent to track and follow a human target with a specific description in dynamic and crowded environments, e.g., “Follow the man in the blue t-shirt.”. The agent must recognize the appearance of the human, follow the correct person described in the instructions, predict their movement trajectory, and keep an appropriate distance while avoiding obstacles [20]. Based on the Habitat 3.0 social navigation environment [56], we collect 544k human-following navigation samples.
The data statistics can be found in Figure 3. Note that all data were collected from synthetic environments, including Habitat-Matterport and Matterport 3D, totaling 861 scenes. We use the default settings of each environment, with a height range from $0.88$ m to $1.25$ m and a robot radius between $0.1$ m and $0.6$ m. This helps avoid overfitting to a specific robot embodiment.
Joint training on synthetic and real-world data. Although we collect navigation data from various environments, the diversity in both observations and instructions remains limited to a certain set of synthetic environments. To incorporate open-world knowledge, we follow previous VLA methods [95, 8], fine-tune Uni-NaVid using open-world video question-answering datasets. We collect 2.3M video question-answering data from publicly available datasets [7, 16, 44], leading to a total of 5.9M training samples. We adopt a two-stage training process: train the visual projector first, and then jointly train both the projector and the large language model. Further training details are provided in the Supplementary Materials.
3.3 Implementation Details
Training configuration. Uni-NaVid is trained on a cluster server with 40 NVIDIA H800 GPUs for approximately 35 hours, totaling 1400 GPU hours. For video data, we sample frames at 1 FPS to remove redundant information between consecutive frames. During training, EVA-CLIP [70] and Vicuna-7B [17] are pre-loaded with default pre-trained weight. Following the training strategy of VLM [46], we optimize the trainable parameters for only 1 epoch.
Benchmark evaluation and real-world deployment. For each navigation task, we adhere to its default settings [37, 64, 18, 31]. Specifically, for EQA [18], the agent executes actions based on the policy until a stop command is output. We then remove the navigation special token <NAV> and query the questions using the navigation history. For real-world deployment, we use a remote server with an NVIDIA A100 GPU to run Uni-NaVid, which processes observations (along with text instructions) and sends commands to a local robot to execute the predicted actions. Uni-NaVid requires approximately 0.2 seconds to generate the next four actions. During navigation, the robot asynchronously uploads the latest observations to the model while executing and pending future actions. Refer to the supplementary video for real-world navigation performance.
4 Experiment
We conduct experiments to evaluate Uni-NaVid on three specific aspects: (1) How does Uni-NaVid perform on individual tasks? (2) Does learning multiple navigation tasks lead to synergistic improvements? (3) Is the key design of our method effective? Due to space limitations, the complete set of protocols, additional results, and visualizations are provided in the Supplementary Materials.
Benchmarks. We evaluate our method on various benchmarks across different navigation tasks. For vision-and-language navigation tasks, we use Val-Unseen splits of R2R [37] and RxR [39] benchmarks. For object goal navigation, we use the validation split of Habitat Matterport 3D dataset [57]. For the embodied question-answering task, we use validation split of MP3D-EQA benchmark [18]. For the human following task, we leverage the Habitat 3.0 social navigation benchmark [56] and extend it to multi-person environments. For video understanding, we follow the evaluation manner of existing methods [44]. We choose ScanQA [7], MSVD [12], MSRVTT [81], and ActivityNet [9] for video question-answering evaluation.
Metrics. To evaluate navigation performance, we follow the standard evaluation metrics [4], including success rate (SR), oracle success rate (OS), success weighted by path length (SPL) [3], trajectory length (TL), following rate (FR) [56], collision rate (CR) [56] and navigation error from goal (NE). Note that the success criteria change among different navigation tasks, we therefore use the default success criteria of each benchmark. For video understanding evaluation, we employ widely used metrics following existing works [7, 44].
4.1 Individual Task Results
Method Observation VLN-CE R2R Val-Unseen
[…]
HPN+DN [38] 7.62 6.31 40.0 36.0 34.0
CMA [25] 10.90 6.20 52.0 41.0 36.0
VLN
[…]
${}^{*}{\dagger}$ [25]
12.23 5.74 53.0 44.0 39.0
Sim2Sim [36] 10.69 6.07 52.0 43.0 36.0
GridMM [74] 13.36 5.11 61.0 49.0 41.0
[…]
ETPNav [1] 11.99 4.71 65.0 57.0 49.0
InstructNav [51] 7.74 6.89 - 31.0 24.0
AG-CMTP [13] 7.90 39.2 23.1 19.1
R2R-CMTP [13] 7.90 38.0 26.4 22.7
LAW [60] 8.89 6.83 44.0 35.0 31.0
CM2 [21] 11.54 7.02 41.5 34.3 27.6
WS-MGMap [14] 10.00 6.28 47.6 38.9 34.3
ETPNav.FF [75] - 5.95 55.8 44.9 30.4
Seq2Seq [37] 9.30 7.77 37.0 25.0 22.0
CMA [37] 8.64 7.37 40.0 32.0 30.0
NaVid [95] 7.63 5.47 49.1 37.4 35.9
Uni-NaVid 9.71 5.58 53.3 47.0 42.7
[…]
Method Observation VLN-CE RxR Val-Unseen
[…]
LAW [60] 4.01 10.87 21.0 8.0 8.0
CM2 [21] 12.29 8.98 25.3 14.4 9.2
WS-MGMap* [14] 10.80 9.83 29.8 15.0 12.1
ETPNav.FF [75] - 8.79 36.7 25.5 18.1
Seq2Seq* [37] 1.16 11.8 5.02 3.51 3.43
CMA* [37] 5.09 11.7 10.7 4.41 2.47
$A^{2}$av [15] 16.8 6.3
NaVid* [95] 10.59 8.41 34.5 23.8 21.2
Uni-NaVid 15.8 6.24 55.5 48.7 40.9
Table 3: Vision-and-language navigation (RxR). Comparison on VLN-CE RxR [40] Val-Unseen. : only trained on VLN-CE R2R.
Comparison on vision-and-language navigation. We evaluate our method with mainstream baselines on two publicly available benchmarks: VLN-CE R2R [37] and RxR [40]. The results are shown in Table 2 and Table 3. We find that our methods achieve SOTA-level performance on both datasets using only RGB videos as observations. In comparison to NaVid [95], which is also a vision language model that is solely trained on VLN data, our approach demonstrates significant improvements, with a $+25.7\%$ increase in Success Rate (SR) on R2R. For zero-shot methods (InstructNav [51] and A2Nav [15]) that use ChatGPT with only text inputs for visual language navigation (VLN), these approaches often face challenges in transitioning between text prompts and visual information, resulting in less than satisfactory outcomes (almost half the SR of Uni-NaVid).
Method Observation HM3D ObjectNav
[…]
DD-PPO [77] 27.9 14.2
Habitat-Web [58] 57.6 23.8
InstructNav [51] 58.0 20.9
PIRLNav-IL [59] 64.1 27.1
PIRLNav-IL-RL [59] 70.4 34.1
OVRL [84] 62.0 26.8
OVRL-v2 [83] 64.7 28.1
Uni-NaVid 73.7 37.1
Table 4: Object goal navigation. Comparison on Habitat Matterport 3D [57] ObjectNav dataset.
Comparison on object goal navigation. We conduct the experiments on HM3D [57] to compare Uni-NaVid with mainstream methods [77, 58, 59, 84, 83] that also learn from ObjectNav data. The results, shown in Table 4, demonstrate that our approach achieves the best performance. Note that methods not utilizing odometry face challenges as they must rely on implicit memory to retain the historical trajectory. Nevertheless, Uni-NaVid still achieves significant gains in SR (+$4.7\%$) and SPL (+$8.8\%$) compared to previous state-of-the-art methods. Additionally, we believe our method’s ObjectNav performance can be further enhanced by incorporating reinforcement learning techniques, as demonstrated by PIRLNav [59] and Poliformer [90].
Method Action Type MP3D EQA
[…]
NaviLLM [98] 44.5
Uni-NaVid 47.3
EQA(habitat-lab) [18] 46.0
NaviLLM [98] 47.4
Uni-NaVid 54.4
Table 5: Embodied question answering. Comparison on Habitat Matterport3D EQA dataset [18].
Comparison on embodied question answering. The evaluation results on MP3D-EQA [18] are presented in Table 5. Despite navigating in continuous environments (CE), our method outperforms existing approaches (e.g., NaviLLM [98] leverage the same evaluation strategy in Sec. 3.3) that operate within discrete landmark-based environments (DE). Moreover, when provided with the ground truth (GT) navigation trajectory, our method shows a significant improvement, demonstrating its ability to understand navigation history effectively.
Method Observation Human Following Dataset
[…]
PoliFormer [90] $\checkmark$ 2.79 20.35 2.93
PoliFormer [90] $\checkmark$ $\checkmark$ 14.67 37.14 4.29
[…]
IBVS [23] $\checkmark$ $\checkmark$ 46.08 62.64 0.84
[…]
Uni-NaVid $\checkmark$ 61.21 71.93 2.07
[…]
Comparison on human following. We compared our method with two most relative methods PoliFormer [90] and IBVS [23]. Since both methods require a specific human bounding box as input, obtained from an upstream algorithm, we use the bounding box from the open-world object detector GroundingDINO [49] and the ground truth provided by the simulator to evaluate the human following performance of the comparison methods under various setups. As shown in Table 6, Uni-NaVid outperforms the comparison methods on both SR (+$21.0\%$) and FR (+$4.4\%$) while maintaining low CR under any setup, even when they use ground truth bounding boxes as input. This demonstrates that Uni-NaVid can effectively infer instructions and follow the correct human, as well as predict the human’s movement patterns accurately.
Figure 4: Visual results of Uni-NaVid in real-world. We deploy Uni-NaVid across diverse environments to execute instructions in a zero-shot setting. We provide both third-person views and the robot’s trajectory, showing effective multi-task navigation performance.
Method ScanQA
EM $\uparrow$ BLUE-1 $\uparrow$ ROUGE $\uparrow$ METEOR   $\uparrow$ CIDEr   $\uparrow$
V.N.+MCAN [89] 19.71 29.46 30.97 12.07 58.23
S.R.+MCAN [89] 20.56 27.85 30.68 11.97 57.56
3D-LLM(flamingo) [26] 23.2 32.60 34.80 13.5 65.6
NaviLLM [98] 26.27 39.73 40.23 16.56 80.77
BridgeQA [54] 31.29 34.49 43.26 16.51 83.75
Uni-NaVid 28.01 46.85 45.74 19.24 94.72
Table 7: Embodied video question answering. Comparison on ScanQA [7] benchmark.
Method MSVD-QA MSRVTT-QA ActivityNet-QA
[…]
VideoLLaMA [91] 51.6 2.5 29.6 1.8 12.4 1.1
VideoChat [42] 56.3 2.8 45.0 2.5 26.5 2.2
VideoChatGPT [52] 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter [47] 67.5 3.7 57.0 3.2 45.7 3.2
Chat-UniVi [32] 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-VID [44] 69.7 3.7 57.7 3.2 47.4 3.3
VideoChat2 [43] 70.0 3.9 54.1 3.3 49.1 3.3
Video-LLaVA [45] 70.7 3.9 59.2 3.5 45.3 3.3
ST-LLM [48] 74.6 3.9 63.2 3.4 50.9 3.3
Uni-NaVid 69.6 3.9 59.3 3.5 51.4 3.7
Table 8: Video question answering. Comparison with leading methods (all based on Vicuna-7B [17]) on VQA benchamarks.
Comparison on video question answering. We first evaluate our method on ScanQA [7] on Tab. 7. Compared to mainstream baselines, we find that Uni-NaVid archives the best performance on four metrics, including BLUE-1 ($+17.9\%$), ROUGE ($+5.7\%$), METEOR ($+16.2\%$), and CIDEr ($+13.1\%$). This proves the superiority of our methods on spatial scene understanding. Note that the EM metric requires an exact match between the question and answer, which is not well-suited to our method, as it is designed to learn from diverse data and generate flexible responses. We also compare our method on open-ended video question answering benchmark in Tab. 8, where we show comparable performance to SOTA methods.
4.2 Qualitative Results in Real-World
Figure 4 visualizes the results of all tasks in real-world environments. Despite training only on synthetic data, Uni-NaVid demonstrates impressive adaptability across various real-world settings, successfully performing multi-task navigation. It effectively aligns navigation history with given instructions, identifies novel objects outside the training set, and avoids obstacles in densely populated areas. Additionally, it reliably follows human movement across diverse environments. Supplementary materials provide numerical results and further visualizations of challenging scenarios.
Figure 5: Comparsion on data scale and multi-task training.
4.3 Ablation Study
We present a visualization of the training strategy’s performance in Figure 5. This comparison contrasts training on a single navigation task with training across multiple tasks. The results (Figure 5 (a)) illustrate the synergistic benefits of multi-task learning, yielding consistent performance gains across all navigation tasks. Notably, VLN, ObjectNav, and EQA exhibit more substantial improvements, while Following shows comparatively smaller gains. We attribute this to the lower dependence of the following task on historical context. Additionally, we investigate the influence of data scale on performance (Figure 5 (b)), observing that performance improves across all navigation tasks with larger data volumes. However, the incremental gain diminishes (from 3M to 6M samples), potentially due to limitations in the data diversity of simulators.
We conduct experiments to evaluate the effectiveness of the training strategy and visual token designs (Tab. 9). Our results indicate that the absence of <NAV> and VQA data leads to a performance decline across all tasks. Notably, the performance drop is most pronounced in EQA, as the lack of <NAV> special token makes the model misinterpret whether it should answer questions or output actions. Additionally, without VQA data, the agent’s ability to answer questions using visual information is greatly impaired.
From the performance of different memory designs, we find that both short-term and long-term memory visual tokens contribute to performance improvements. In particular, the VLN task shows the most significant performance drop ($-80.3\%$ SR) when visual memory is removed, as the lack of memory hinders the alignment of visual history with instructions. For the Following task, the absence of memory results in only a minor performance decline ($-8\%$ SR), as this task primarily relies on recent frames to track the target. Additional ablation studies are provided in the Supplementary Material.
[…]
No $<$Nav$>$ token 35.2 69.1 20.4 55.1
No VQA data 40.5 50.6 1.19 58.8
Curr. 9.61 44.3 32.5 56.3
Curr.+Short. 39.7 67.8 44.1 59.7
Curr.+Short.+Long. 48.7 73.7 47.3 61.2
Table 9: Ablation study on both training strategy and architecture.
5 Conclusion
We introduce a vision-language-action model, Uni-NaVid, designed to acquire general embodied navigation skills through multi-task navigation data. To efficiently encode the extended video sequences encountered in navigation, we develop an online visual token merging strategy that separately processes current observations, short-term observations, and long-term observations. This design allows our approach to operate at an average speed of 5 Hz. Extensive experiments across publicly available benchmarks reveal our model’s superior performance. The ablation studies demonstrate the synergy achieved by learning across multiple navigation tasks and prove the effectiveness of our key designs. In the future, we will explore the synergy between more fundamental embodied skills, integrating both manipulation and navigation.