### v2

# Uni-NaVid: A Video-based Vision-Language-Action Model for Unifying Embodied Navigation Tasks

Abstract
Embodied Navigation is a fundamental capability for intelligent robots, requiring robots to follow human commands and move autonomously within physical environments. Despite significant advancements, most existing navigation approaches are tailored to specific navigation tasks, such as instruction following, searching objects, answering questions, tracking people, and more. However, the increasing demands on advanced embodied navigation pose the challenge of designing a practical navigation agent that can incorporate multiple navigation tasks naturally and benefits from the synergy between these tasks. To this end, we present Uni-NaVid, a video-based vision-language-action (VLA) model to unify different paradigms of navigation tasks and improve navigation performance by encouraging the synergy among different navigation sub-tasks. This VLA model can directly take natural language instructions and RGB video streams as inputs and output low-level robotic actions in an end-to-end manner. To efficiently process extensive RGB video streams, we propose an online token merge strategy that spatially and temporally consolidates similar visual information which improves the inference speed to 5 Hz. For training Uni-NaVid, we collect 3.6 million navigation data samples across different navigation tasks. Extensive experiments on diverse navigation benchmarks demonstrate that Uni-NaVid achieves state-of-the-art performance within a unified framework by using only ego-centric RGB video as inputs. Additionally, real-world experiments confirm the model’s effectiveness and efficiency, shedding light on its strong generalizability.
Figure 1: Uni-NaVid learns general navigation skills across four embodied navigation tasks using 3.6 million navigation samples. Uni-NaVid only takes online RGB video frames and language instructions as input and output actions, achieving general navigation ability in a real-world deployment.
00footnotetext: Author (e-mail: zhngjizh@gmail.com). $\dagger$ Corresponding authors (e-mail: zhangzz@galbot.com, hewang@pku.edu.cn).
I Introduction
Embodied navigation [112, 85] is a critical capability for intelligent robots and has drawn significant attention in the robotics community. For successful embodied navigation, robots must be able to move autonomously within physical environments based on human instructions. However, navigation tasks vary significantly, and most existing studies are designed for specific tasks, e.g., vision-and-language navigation [44, 46], object goal navigation [12], embodied question answering [21, 90], and following [113, 36, 68]. Consequently, most current approaches are developed to address only one type of navigation task, often relying on specialized modules and task-specific datasets. This narrow scope limits their applicability to multi-purpose navigation applications and prevents these methods from leveraging potential synergies across diverse navigation tasks.
Developing a versatile navigation model presents significant challenges, as it requires the unification of navigation task modeling and the integration of heterogeneous data for joint use. Initial efforts adopt imitation learning (IL) [85, 93, 66] or reinforcement learning (RL) [106, 97] to learn general navigation skills in simulation environments or limited diverse real-world environments. However, due to the limited rendering quality and diversity of simulators, these approaches often encounter the “sim-to-real” gap and suffer from poor generalization across diverse navigation tasks [27, 5, 40]. Recent studies [120, 114, 61, 60, 78] have attempted to achieve a higher degree of unification using pre-trained large language models (LLMs). However, due to the low frequency of LLM inference, they simplify the problem to some extent by adopting discretized modeling approaches. They rely on pre-defined graphs for decision-making learning, which sacrifices output flexibility and introduces additional challenges for real-world deployment.
In this work, we propose Uni-NaVid, a video-based Vision-Language-Action (VLA) model for unifying diverse commonly demanded navigation tasks (Tab. I). Uni-NaVid  takes egocentric RGB video streams and natural language instructions as inputs, and directly generates low-level actions for navigation in continuous environments. To achieve multi-task navigation while supporting efficient navigation, Uni-NaVid extend video-based VLM [51] by incoprating two key components: (1) an efficient VLA architecture based on an online token merge mechanism, which enables efficient processing of online-captured video streams for LLM inference; and (2) an extensive collection of 3.6M samples across four widely studied navigation tasks. We provide a detailed elaboration below:
During navigation, the agent is required to process a substantial volume of online captured frames, which results in memory overload and computational latency, particularly in LLM-based approaches [111, 61]. To this end, we propose an online token merging mechanism to compress near historical frames with a relatively low ratio while compressing far historical frames with a relatively high ratio. This merging mechanism operates in an on-the-fly manner, maximizing the reuse of previous navigation history. In this way, Uni-NaVid learn compact representations that maintain not only fine-grained spatial information but also structured temporal information, thus speeding up the model inference by reducing the token number. Besides, Uni-NaVid adopts a foresight prediction to generate actions for a future horizon at once instead of step-by-step. This enables Uni-NaVid to achieve 5Hz inference, facilitating the deployment of a non-blocking navigation robot powered by a VLA model in real-world environments (Please refer to the supplementary video).
We aim to build  Uni-NaVid as a versatile multi-task navigation agent, incorporating four widely demanded navigation tasks: vision-and-language navigation, object-goal navigation, embodied question answering, and human following. These tasks are distinct from each other, with varying task settings and objectives. Specifically, for the human-following task, we construct a new language-guided human-following benchmark for data collection and evaluation. Finally, we collect 3.6M navigation samples based on diverse navigation tasks with different simulation environments. Additionally, inspired by the success of manipulation VLAs [9], we further integrate 2.3M real-world internet data samples for Video Question Answering (VQA) [7, 51] and video captioning [19] as auxiliary tasks. This integration aims to enhance scene understanding and promote sim-to-real generalization.
We conduct extensive experiments on benchmarks across the aforementioned four navigation tasks and compared our method with strong baselines specifically designed for each task. Utilizing only RGB video streams and instructions as inputs, our method demonstrates the superiority of a single VLA model across diverse benchmarks, achieving SOTA or SOTA-comparable performance. Furthermore, comprehensive ablation studies validate the synergistic benefits of learning multiple navigation tasks jointly. Finally, real-world experiments demonstrate that Uni-NaVid achieves non-blocking navigation exhibiting impressive robustness in handling diverse instructions and environments. We believe our work serves merely as a starting point for general-purpose navigation, and we will release the full source code to benefit the community.
II Related Works
Figure 2: Pipeline of Uni-NaVid. Our method takes only single-view RGB frames $\{\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\}$ and a natural language instruction $\mathcal{I}$ as input. For each frame, we extract 64 visual tokens using the vision encoder and then use online token merging to accelerate the model while retaining compact visual information. The merged tokens and instruction tokens are sent to the large language model to obtain actions for navigation or answers for embodied question-answering.
Methods Action Embodied Navigation Tasks
D.E. C.E. VLN [44] ObjNav [76] EQA [90] Follow [68]
VLMaps [34]
NaviLLM [114]
InstructNav [61]
Poliformer [106]
Uni-NaVid
TABLE I: Task and setting comparison. Uni-NaVid is developed to address four embodied navigation tasks, generating action outputs in continuous environments. C.E.: Continuous Environment; D.E.: Discrete Environment.
Multi-Task Embodied Navigation. Embodied navigation [2, 94, 112] requires agents to navigate in unseen environments based on human instructions. There is extensive literature on embodied navigation; here, we focus on four mainstream tasks that involve both visual information and language instructions: Vision-and-Language Navigation [4, 44, 46], Object Goal Navigation [12], Embodied Question Answering [21], and Human Following [37, 68, 118, 119]. Early efforts [85, 93, 66, 97] towards a generalist-embodied navigation model involved multi-task navigation datasets and directly learning navigation skills, showing initial success in multi-task performance. However, these methods experienced performance drops when deployed in novel environments, especially in real-world settings. In recent years, advanced approaches [114, 61, 35, 60, 121, 78] have leveraged the generalization capabilities of large language models to improve multi-task navigation. These models show promising generalizability across navigation tasks but rely on extensive prompting, which impacts time efficiency. In contrast, our video-based large language model is trained end-to-end for multi-task navigation, offering robust generalization and computational efficiency for tasks like human following.
Embodied Navigation Datasets. To train and evaluate the performance of a policy for embodied navigation tasks, a large body of datasets and corresponding benchmarks have been proposed [23, 123, 59, 64]. These datasets play a crucial role in the embodied navigation community. Here, we review the datasets most relevant to our methods. For vision-and-language navigation, the most widely used datasets are Room-2-Room (R2R) [4] and Room-cross-Room (RxR) [47], which provide navigation instructions and ground truth trajectories of landmarks. We focus on a variant of R2R and RxR in continuous environments, called VLN-CE [44], which is more practical for real-world applications. For object goal navigation, there are several famous benchmarks such as HM3D [69], MP3D [11], and Aithor [124], which are built on various scene environments and simulators. Here, we leverages the HM3D dataset on Habitat [76], which shares the same action settings as VLN-CE. For embodied question answering (EQA), there are diverse datasets focusing on different attributes of EQA, such as MP3D-EQA [90], MT-EQA [103], Graph-EQA [83], and MX-EQA [38]. We select MP3D-EQA, which is well-maintained with the latest baselines. For human-following [116, 117] benchmarks, there is currently no benchmark that provides textual descriptions of humans. Therefore, we have self-built a textual description-based human-following benchmark using Habitat 3.0 [68]. Note that new benchmarks are consistently being proposed, covering a diverse range of navigation attributes. However, our goal is to train and evaluate our method on mainstream datasets to clearly justify the performance of our approach.
Large Language Models for Navigation. Large Language Models (LLMs)[20, 54, 122] have been introduced into robotic navigation due to their generalization capabilities in understanding and planning. One straightforward approach[120, 61, 60, 78] is to use off-the-shelf large language models in a zero-shot manner. These methods employ visual foundation models [22, 54] to describe surrounding environments in text format, prompting the language model to select landmarks that guide the agent. However, abstracting dense visual information into text and relying on discrete landmarks results in sparse environmental observations and is limited to static environments. Another approach [111, 106] trains a video-based large language model end-to-end with low-level actions to enable continuous movement. However, it faces efficiency challenges in long-horizon tasks. In contrast, Uni-NaVid implements an online visual token merging strategy, optimizing training efficiency for long-horizon tasks and supporting non-blocking execution in real-world environments.
III Problem Formulation
[…]
Overview. As illustrated in Figure 2, Uni-NaVid  is composed of three main components: a vision encoder, an online token merge mechanism and a large language model (LLM). First, the online captured video stream is encoded by the vision encoder (EVA-CLIP [82] in implementation) to extract frame-wise visual features in the form of tokens, which we denote them as visual tokens. The visual tokens are then spatially and temporally merged by leveraging an online token merge mechanism. Next, the merged visual tokens are projected with an MLP projector into a feature space aligned with language tokens, which are referred to as visual observation tokens. As common, the instructions are also tokenized as a set of tokens, known as language observation tokens. Both the visual observation tokens and language observation tokens are concatenated and passed to the Large Language Model (LLM), which infers four action tokens that represent the next four actions.
IV Model of Uni-NaVid
IV-A Observation Encoding.
Given the ego-centric video up to time $T$, denoted by $\mathcal{O}_{T}=\{\mathbf{x}_{1}.\cdots,\mathbf{x}_{T}\}$, we encode the video to a sequence of visual features in the form of tokens. For each frame $\mathbf{x}_{t}$, we first get its visual feature tokens $\mathbf{X}_{t}\in\mathbb{R}^{N_{x}\times C}$ with a vision encoder (EVA-CLIP [82] in implementation), where $N_{x}$ is the patch number ($N_{x}$ is set to 256) and $C$ is the embedding dimension.
$\mathbf{X}_{1:T}=Encoder(\mathbf{x}_{1:T})$ (1)
The visual features provide rich information that enables the agent to understand its navigation history and plan subsequent actions. However, during navigation, the progressively increasing number of visual tokens $(T\times N_{x})$ results in progressively longer inference times for the LLM (typically 1–2 seconds per inference) [111]. This increased latency renders LLM-based navigation impractical for deployment in real-world environments.
IV-B Online Visual Token Merging
To reduce the number of visual tokens while preserving sufficient navigation visual information, we design an token merging mechanism. This strategy is based on the key insight that recent observations are more critical for navigation, and that visual information between consecutive frames (temporally) and within neighboring pixels (spatially) may be redundant.
Visual token grouping. Drawing inspiration from the Atkinson-Shiffrin memory model [6, 80], we categorize visual tokens into current visual tokens $\mathbf{X}_{\text{curr}}$, short-term visual tokens $\mathbf{X}_{\text{short}}$, and long-term visual tokens $\mathbf{X}_{\text{long}}$. These visual tokens are grouped based on their timestamps relative to the current frame $T$ and for each group of visual tokens, we apply a grid pooling operation at different pooling resolutions:
$\mathbf{X}_{1:T}=\begin{cases}\mathbf{X}_{\text{curr}}=GridPool(\mathbf{X}_{t},\alpha_{\text{curr}}),&\text{if t = T}\\ \mathbf{X}_{\text{short}}=GridPool(\mathbf{X}_{t},\alpha_{\text{short}}),&\text{if t $\in$ [T-B, T)}\\ \mathbf{X}_{\text{long}}=GridPool(\mathbf{X}_{t},\alpha_{\text{long}}),&\text{if t $\in$ [1, T-B)}\end{cases}$ (2)
where GridPool(·) is a grid pooling operation  [51, 111], spatially squeezing the tokens from $N_{x}$ to $\frac{N_{x}}{\alpha^{2}}$, and $B$ (set to $64$) is the length of the buffer of shorter memory. Here, we adopt the $\alpha_{\text{curr}}=2$, $\alpha_{\text{short}}=8$, $\alpha_{\text{long}}=16$, leads to visual tokens as $\mathbf{X}_{\text{curr}}\in\mathbb{R}^{64\times C}$, $\mathbf{X}_{\text{short}}\in\mathbb{R}^{4\times C}$, $\mathbf{X}_{\text{long}}\in\mathbb{R}^{1\times C}$, respectively. Here, current visual tokens $\mathbf{X}_{\text{curr}}$ encapsulate comprehensive visual information, enabling the agent to perceive its immediate environment and plan subsequent trajectories. Meanwhile, $\mathbf{X}_{\text{short}}$ and $\mathbf{X}_{\text{long}}$ capture temporally rich information from the captured video stream, facilitating the agent’s comprehension of its navigation history.
It should be noted that these hyperparameters are obtained through empirical experimentation to achieve an optimal balance between manageable token numbers and adequate visual information representation. These hyperparameters can be further adjusted when memory capacity and computational resources are not limiting factors. We provide a detailed explanation and ablation study of $\alpha$ in the supplemental material.
Online visual token process. During the navigation process, the agent consistently observes new frames. However, performing encoding and grouping (Eq. 2) for all frames at each step would be computationally intensive. To address this, we implement an online visual token processing mechanism that maximizes the reuse of previously generated visual tokens. Specifically, when a new frame at time $T+1$ is received, we apply grid pooling exclusively to the most recent visual tokens at time $T$ and the oldest short-term visual tokens at time $T-B$. These processed tokens are then integrated into the short-term and long-term visual tokens, respectively:
$\mathbf{X}_{\text{curr}\rightarrow\text{short}}$ $=GridPool(\mathbf{X}_{\text{curr}},\frac{\alpha_{\text{short}}}{\alpha_{\text{curr}}}),$ (3)
$\mathbf{X}_{\text{short}\rightarrow\text{long}}$ $=GridPool(\mathbf{X}_{\text{short}},\frac{\alpha_{\text{long}}}{\alpha_{\text{short}}}).$ (4)
To prevent the linear growth of long-term visual tokens $\mathbf{X}_{\text{Long}}$, we further perform token merging on the long-term visual tokens by combining adjacent tokens that exhibit high similarity, following the approach of VLM-based methods [8, 80]. Specifically, we merge the long-term visual tokens based on the cosine similarity between $\mathbf{X}_{\text{short}\rightarrow\text{long}}$ and the most recent long-term visual tokens $\mathbf{X}_{\text{long}}$ at time $T-B-1$. If the similarity exceeds a predefined threshold $\tau$, we merge them according to the number of frames previously merged (denoted as $K$) in the latest long-term visual tokens:
$\mathbf{X}_{\text{long}}$ $=\frac{1}{K+1}\left(K\mathbf{X}_{\text{long}}+\mathbf{X}_{\text{short}\to\text{long}}\right),$ (5)
subject to $\cos\left(\mathbf{X}_{\text{long}},\mathbf{X}_{\text{short}\to\text{long}}\right)>\tau.$ (6)
We insert new long-term visual tokens $\mathbf{X}_{\text{short}\to\text{long}}$ when their similarity falls below a threshold $\tau$ (empirically set to $\tau=0.95$ [80]), indicating that they contain relatively distinct visual information. This online visual token processing preserves the navigation visual history in a highly compact form (with a length of $M\ll T-B-1$). Notably, only visual tokens at the boundaries of groups require parallelizable grid pooling, making the process computationally efficient and naturally suited for online deployment in real-world navigation tasks. We give a description of our token merging technique at Algorithmn 1.
Algorithm 1 Online Visual Token Merging
1:
2:
  • Total number of frames $T$
  • Short memory buffer length $B$
  • Grid pooling scales: $\alpha_{\text{curr}}$, $\alpha_{\text{short}}$, $\alpha_{\text{long}}$
  • Current visual tokens: $\mathbf{X}_{T}\in\mathbb{R}^{N_{x}\times C}$
  • Previously merged tokens: $\mathbf{X}_{\text{curr}}$, $\mathbf{X}_{\text{short}}$, $\mathbf{X}_{\text{long}}$
  • Number of frames merged in the last tokens of long memory: $K$
3:
4:
  • Updated merged tokens: $\mathbf{X}^{\prime}_{\text{curr}}$, $\mathbf{X}^{\prime}_{\text{short}}$, $\mathbf{X}^{\prime}_{\text{long}}$
  • Updated number of frames merged in the last tokens of long memory: $K^{\prime}$
5:
6:if $T==1$ then $\triangleright$ First frame, empty history tokens
7:     $\mathbf{X}^{\prime}_{\text{short}},\mathbf{X}^{\prime}_{\text{long}}\leftarrow[]$
8$\triangleright$ Update short-term visual tokens
9:     $\mathbf{X}_{\text{curr}\rightarrow\text{short}}\leftarrow GridPool(\mathbf{X}_{\text{curr}},\frac{\alpha_{\text{short}}}{\alpha_{\text{curr}}})$
10:     $\mathbf{X}^{\prime}_{\text{short}}\leftarrow\mathbf{X}_{\text{short}}+[\mathbf{X}_{\text{curr}\rightarrow\text{short}}]$
11:end if
12:$\mathbf{X}^{\prime}_{\text{curr}}\leftarrow GridPool(\mathbf{X}_{T},\alpha_{\text{curr}})$triangleright New current visual token
13:if $T>B+1$ then $\triangleright$ Out of short-term tokens buffer
14:     $\mathbf{X}_{\text{short}\rightarrow\text{long}}\leftarrow GridPool(\mathbf{X}_{\text{short}}[0],\frac{\alpha_{\text{long}}}{\alpha_{\text{short}}})$
15:     $\mathbf{X}^{\prime}_{\text{short}}\leftarrow\mathbf{X}_{\text{short}}[1:]$
16:     $s\leftarrow\cos(\mathbf{X}_{\text{long}}[-1],\mathbf{X}_{\text{short}\rightarrow\text{long}})$
17:     if $T>B+2\ \text{and}\ s>\tau$ then $\triangleright$ Fuse long-term tokens
18:         $\mathbf{X}_{\text{last\_long}}\leftarrow\frac{1}{K+1}(K\mathbf{X}_{\text{long}}[-1]+\mathbf{X}_{\text{short}\rightarrow\text{long}})$
19:         $\mathbf{X}^{\prime}_{\text{long}}\leftarrow\mathbf{X}_{\text{long}}[:-1]+[\mathbf{X}_{\text{last\_long}}]$
20:         $K^{\prime}\leftarrow K+1$
21:     $\triangleright$ Add new long-term token
22:         $\mathbf{X}^{\prime}_{\text{long}}\leftarrow\mathbf{X}_{\text{long}}+[\mathbf{X}_{\text{short}\rightarrow\text{long}}]$
23:         $K^{\prime}\leftarrow 1$
24:     end if
25:end if
Compared to existing video-based large language models [111, 80, 51], this online merging strategy significantly reduces inference time, achieving an average of 0.2 seconds per inference. This improvement becomes increasingly notable when handling longer video sequences. A detailed analysis of time efficiency is provided in the Supplementary Materials.
IV-C Action Planning
After obtaining the merged visual tokens from semantic features  [82], we adopt established practices in Vision-and-Language models  [54, 51] to perform vision-language alignment, enabling the large language model (LLM) to effectively interpret visual information. Specifically, we leverage a cross-modality projector $P_{V}(\cdot)$ to project all merged visual tokens $X_{\text{merged}}=\{\mathbf{X}_{\text{long}},\mathbf{X}_{\text{short}},\mathbf{X}_{\text{curr}}\}$ into visual observation tokens that are compatible with the LLM’s input representation space:
$\mathbf{E}^{V}_{T}=P_{V}(\mathbf{X}_{\text{merged}}),$ (7)
[…]
Input: $\{Long\_term\_tokens\}\{Shot\_term\_tokens\}\\ \{Current\_tokens\}<\!NAV\!>\{Instruction\}$ Output: $<\!Action\_0\!>\!<\!Action\_1\!>\!<\!Action\_2\!>\!\\ \!<\!Action\_3\!>$
[…]
V Data Collection and Training
To train  Uni-NaVidfor mastering multi-navigation tasks, it is crucial to gather extensive and diverse navigation data across various tasks and environments. However, directly collecting large amounts of real-world navigation data can be prohibitively expensive. To address this challenge, we propose two key strategies for training Uni-NaVid: First, we collect multi-task navigation data from a wide range of synthetic environments (totaling 861 scenes) using a uniform input and output format, enabling Uni-NaVidto acquire general navigation skills. Second, we co-tune Uni-NaVidwith real-world video-based question-answering data, enhancing its ability to interpret real-world images and supporting its open-vocabulary knowledge acquisition.
Figure 3: Visualization of training data. We visualize the combination of training data (5.9M), video frame counts, and the most common words in navigation instructions.
V-A Multi-Task Navigation Data.
We collect the largest multi-task navigation dataset to date within the Habitat simulator environment [76], comprising 3.6 million samples across four distinct navigation tasks, as described below. All tasks are curated within a unified framework. A detailed data collection strategy is provided in the Supplementary Materials.
(A) Vision-and-language navigation [44, 46] require the agent to interpret and ground instructions in visual observations, effectively combining linguistic and visual information to make sequential decisions. Specifically, the agent has to navigate based on landmarks and motions described in the text and stop nearby the correct destination. Here, we collect 2.4M navigation samples of mainstream VLN datasets, VLN-CE R2R [44] and RxR [47], that focus on continuous environments.
(B) Object Goal Navigation [76] involves an agent navigating an environment to locate a specific object based on provided visual or linguistic cues. This task evaluates the agent’s ability to perceive objects, understand scene layout, and execute efficient search strategies. We collected 483k samples from datasets in the Habitat Matterport 3D dataset (HM3D ObjectNav) [70]. Note that, in HM3D ObjectNav, the agent is required to locate objects from a predefined category set (e.g., sofa, chair, and bed). Nevertheless, experiments demonstrate that our method generalizes to SOTA-level open-vocabulary object goal searching, as shown in Table V.
(C) Embodied question answering [90] requires the agent to navigate to the related area for question answering. It involves spatial reasoning, object description, and understanding contextual information, requiring the ability to integrate perception, language comprehension, and decision-making. Following the setup in main stream EQA methods [21, 90], the agent first navigates to the target related to the question, issues a stop action, and then provides an answer. We collect 240k video-action samples and 10k video-answering samples on the MP3D-EQA dataset [21] on Matterport 3D environments [11].
(D) Human following [37, 25] requires the agent to track and follow a human target with a specific description in dynamic and crowded environments, e.g., “Follow the man in the blue t-shirt.”. The agent must recognize the appearance of the human, follow the correct person described in the instructions, predict their movement trajectory, and keep an appropriate distance while avoiding obstacles.
Figure 4: Language-described human following benchmark. We construct our human-following benchmark based on Habitat 3.0 [68] by incorporating textual descriptions for each avatar (eight in total, top row). The robot is required to comprehend these descriptions and accurately follow the designated individual in crowded environments.
However, there is currently no human-following dataset that supports language-described human following in crowded environments (multi-person scenarios). To this end, we extend the Habitat 3.0 social navigation benchmark [68] by (1) adding textual descriptions for each avatar (8 in total, as illustrated in Fig. 4), (2) introducing additional distracting human avatars to simulate challenging real-world environments, and (3) deploying the robot and humans in the Habitat Matterport 3D dataset [101], which offers photo-realistic rendering quality and diverse large-scale scenes. The robot and target human are initialized nearby (using the same setting as  [68]), with randomly moving distracting human avatars. Based on this setup, we collected 544k human-following navigation samples. We also add a detailed description in Supplementary Material. This benchmark will also be released to benefit the navigation community.
Unified navigation samples. The data statistics are presented in Figure 3. It is worth noting that the number of samples in VLN is relatively larger compared to other tasks. This is because VLN [44, 46] requires the agent to navigate all landmarks described in the instructions, which often results in longer trajectories and, consequently more video-action samples. Here, we collect all navigation samples in a uniform format, including an egocentric RGB video, a natural language instruction, and four corresponding future actions. All data were collected from synthetic scenes across the Habitat-Matterport 3D (HM3D) and Matterport 3D (MP3D) datasets. We use the default settings of each environment, with a height range of $0.88$ m to $1.25$ m and a robot radius between $0.1$ m and $0.6$ m. This approach helps prevent overfitting to a specific robot embodiment. This approach helps prevent overfitting to a specific robot embodiment. Note that while there exist insightful techniques [24, 29] investigating navigation for robots of general sizes, our focus is primarily on uniform multi-task navigation.
V-B Training Strategy of Uni-NaVid
[…]
Training configuration. Uni-NaVid is trained on a cluster server with 40 NVIDIA H800 GPUs for approximately 35 hours, totaling 1400 GPU hours. For video data, we sample frames at 1 FPS to remove redundant information between consecutive frames. During training, the vision encoder (EVA-CLIP [82]) and large language model (Vicuna-7B [20]) are pre-loaded with default pre-trained weight. Following the training strategy of VLM [54], we optimize the trainable parameters for only 1 epoch.
VI Experiment
We conduct experiments to evaluate Uni-NaVid on three specific aspects: (1) How does Uni-NaVid perform on individual tasks? (2) Does learning multiple navigation tasks lead to synergistic improvements? (3) Is the key design of our method effective? To evaluate the general-purpose navigation method, we conduct extensive experiments on individual navigation tasks, employing corresponding strong baselines. Additional details are provided in the supplemental material.
Benchmarks. We evaluate our method on various benchmarks across different navigation tasks. Given the diversity of benchmarks spanning various environments and simulators, we meticulously verify the scene splits to ensure no overlap exists between the training and validation scenes across benchmarks.
  • Vision-and-language navigation: We test our method on the validation splits of the VLN-CE R2R [44] and RxR [46] benchmarks.
  • Object goal navigation: We use the validation split of the Habitat Matterport 3D (HM3D) dataset [70], which requires the agent to find target objects from six categories (sofa, chair, TV, bed, toilet, and plant) in unseen environments. Moreover, to test generalizability, we also evaluate our method on the HM3D-OVON dataset [101], an open-vocabulary object navigation benchmark, in a zero-shot manner.
  • Embodied question-answering: We use the validation split of the MP3D-EQA benchmark [90]. Additionally, we conduct experiments on the more recent Embodied Video Question Answering benchmark, OpenEQA [63].
  • Human following: We evaluate our method alongside mainstream approaches on our proposed language-described human following benchmark.
  • Video understanding: We follow the evaluation procedures of existing VQA methods [51]. We choose the ScanQA [7], MSVD [13], MSRVTT [96], and ActivityNet [10] datasets.
Metrics. To evaluate navigation performance, we follow the standard evaluation metrics [4], including success rate (SR), oracle success rate (OS), success weighted by path length (SPL) [3], trajectory length (TL), following rate (FR) [68], collision rate (CR) [68] and navigation error from goal (NE). Note that the success criteria change among different navigation tasks, we therefore use the default success criteria of each benchmark. For video understanding evaluation, we employ widely used metrics following existing works [7, 51].
VI-A Deployment Details of Uni-Navid.
Benchmark evaluation. For each navigation task, we adhere to the default settings of each navigation task [44, 76, 21, 37]. All tasks take an online captured RGB video (capturing one frame after each action) and a textual instruction as inputs, and output the next four actions (Sec. IV-C). The robot then executes the predicted actions and calls STOP once the first predicted action is a stop action. For VLN and EQA tasks, we directly use the text instruction provided by the benchmark episodes. For human following and object goal navigation, we transform the target information into an instruction by adding prefixes such as ”Search for” or ”Follow.” Further details can be found in the supplemental material.
It is worth noting that for EQA [21] task, the agent executes navigation actions until a stop command is issued. We then remove the navigation-specific token <NAV> and query the questions using the navigation history. This strategy alleviates the ambiguity for the LLM in deciding whether to navigate or answer a question (See Table X).
Real-world deployment. For real-world deployment, we utilize a remote server with an NVIDIA A100 GPU to run Uni-NaVid, which processes observations (along with text instructions) and sends commands to a local robot to execute the predicted actions. Uni-NaVid requires approximately 0.2 seconds to generate the next four actions. During navigation, the robot asynchronously compresses and uploads the latest observations to the model while executing pending actions. Refer to the supplementary video for real-world navigation performance.
VI-B Individual Task Results
Method Observation VLN-CE R2R Val-Unseen
[…]
HPN+DN [45] 7.62 6.31 40.0 36.0 34.0
CMA [31] 10.90 6.20 52.0 41.0 36.0
VLN
[…]
${}^{*}{\dagger}$ [31]
12.23 5.74 53.0 44.0 39.0
Sim2Sim [43] 10.69 6.07 52.0 43.0 36.0
GridMM [87] 13.36 5.11 61.0 49.0 41.0
[…]
ETPNav [1] 11.99 4.71 65.0 57.0 49.0
InstructNav [61] 7.74 6.89 - 31.0 24.0
AG-CMTP [15] 7.90 39.2 23.1 19.1
R2R-CMTP [15] 7.90 38.0 26.4 22.7
LAW [73] 8.89 6.83 44.0 35.0 31.0
CM2 [26] 11.54 7.02 41.5 34.3 27.6
WS-MGMap [16] 10.00 6.28 47.6 38.9 34.3
ETPNav.FF [88] - 5.95 55.8 44.9 30.4
Seq2Seq [44] 9.30 7.77 37.0 25.0 22.0
CMA [44] 8.64 7.37 40.0 32.0 30.0
NaVid [111] 7.63 5.47 49.1 37.4 35.9
Uni-NaVid 9.71 5.58 53.3 47.0 42.7
[…]
Method Observation VLN-CE RxR Val-Unseen
[…]
LAW* [73] 4.01 10.87 21.0 8.0 8.0
CM2* [26] 12.29 8.98 25.3 14.4 9.2
WS-MGMap* [16] 10.80 9.83 29.8 15.0 12.1
ETPNav.FF [88] - 8.79 36.7 25.5 18.1
Seq2Seq* [44] 1.16 11.8 5.02 3.51 3.43
CMA* [44] 5.09 11.7 10.7 4.41 2.47
$A^{2}$av [17] 16.8 6.3
NaVid* [111] 10.59 8.41 34.5 23.8 21.2
Uni-NaVid 15.8 6.24 55.5 48.7 40.9
TABLE III: Vision-and-language navigation (RxR). Comparison on VLN-CE RxR [47] Val-Unseen. : only trained on VLN-CE R2R.
[…]
Method Observation HM3D ObjectNav
[…]
DD-PPO [91] 27.9 14.2
Habitat-Web [71] 57.6 23.8
InstructNav [61] 58.0 20.9
PIRLNav-IL [72] 64.1 27.1
PIRLNav-IL-RL [72] 70.4 34.1
OVRL [99] 62.0 26.8
OVRL-v2 [98] 64.7 28.1
Uni-NaVid 73.7 37.1
TABLE IV: Object goal navigation. Comparison on Habitat Matterport 3D [70] ObjectNav dataset.
Comparison on object goal navigation. We conduct the experiments on HM3D [70] to compare Uni-NaVid with mainstream methods [91, 71, 72, 99, 98] that also learn from ObjectNav data. The results, shown in Table IV, demonstrate that our approach achieves the best performance. Note that methods not utilizing odometry face challenges as they must rely on implicit memory to retain the historical trajectory. Nevertheless, Uni-NaVid still achieves significant gains in SR (+$4.7\%$) and SPL (+$8.8\%$) compared to previous state-of-the-art methods. Additionally, we believe our method’s ObjectNav performance can be further enhanced by incorporating reinforcement learning techniques, as demonstrated by PIRLNav [72] and Poliformer [106].
Method VAL SEEN VAL SEEN SYNONYMS VAL UNSEEN
[…]
BC 11.1 4.5 9.9 3.8 5.4 1.9
DAgger 11.1 4.5 9.9 3.8 5.4 1.9
RL 18.1 9.4 15.0 7.4 10.2 4.7
BCRL 39.2 18.7 27.8 11.7 18.6 7.5
DAgRL 41.3 21.2 29.4 14.4 18.3 7.9
VLFM [100] 35.2 18.6 32.4 17.3 35.2 19.6
DAgRL+OD [101] 38.5 21.1 39.0 21.4 37.1 19.8
Uni-NaVid 41.3 21.1 43.9 21.8 39.5 19.8
TABLE V: Object goal navigation. Comparison on HM3D-OVON [101]. ${}^{*}:$ denotes zero-shot methods.
To evaluate the generalization ability for open-vocabulary objects, we evaluate our method on the open-vocabulary object goal navigation benchmark (HM3D-OVON [101]) in a zero-shot manner. The results in Table V demonstrate that our method achieves significant improvement over the zero-shot method (VLFM [100]) and even outperforms the fine-tuned method (DAgRL+OD [101]) on the VAL SEEN and VAL UNSEEN splits. This proves the generalizability of our method.
Method Action Type MP3D EQA
[…]
NaviLLM [114] 44.5
Uni-NaVid 47.3
EQA(habitat-lab) [21] 46.0
NaviLLM [114] 47.4
Uni-NaVid 54.4
TABLE VI: Embodied question answering. Comparison on Habitat Matterport3D EQA dataset [21].
Comparison on embodied question answering. The evaluation results on MP3D-EQA [90] are presented in Table VI. Despite navigating in continuous environments (CE), our method outperforms existing approaches (e.g., NaviLLM [114] leverage the same evaluation strategy in Sec. VI-A) that operate within discrete landmark-based environments (DE). Moreover, when provided with the ground truth (GT) navigation trajectory, our method shows a significant improvement, demonstrating its ability to understand navigation history effectively. We also report our performance on the more challenging EM-EQA benchmark, OpenEQA [63], in the Supplemental Material, which includes more complex questions. Our method achieves comparable performance to GPT-4V with scene captions [63].
Method Observation Human Following Dataset
[…]
PoliFormer [106] $\checkmark$ 2.79 20.35 2.93
PoliFormer [106] $\checkmark$ $\checkmark$ 14.67 37.14 4.29
[…]
IBVS [28] $\checkmark$ $\checkmark$ 46.08 62.64 0.84
[…]
Uni-NaVid $\checkmark$ 61.21 71.93 2.07
[…]
Comparison on human following. We compared our method with two most relative methods PoliFormer [106] and IBVS [28]. Since both methods require a specific human bounding box as input, obtained from an upstream algorithm, we use the bounding box from the open-world object detector GroundingDINO [58] and the ground truth provided by the simulator to evaluate the human following performance of the comparison methods under various setups. As shown in Table VII, Uni-NaVid outperforms the comparison methods on both SR (+$21.0\%$) and FR (+$4.4\%$) while maintaining low CR under any setup, even when they use ground truth bounding boxes as input. This demonstrates that Uni-NaVid can effectively infer instructions and follow the correct human, as well as predict the human’s movement patterns accurately. We include additional human-following experiments in various environments, such as HSSD [42] and MP3D [11], in the Supplemental Material. Our method consistently demonstrates SOTA performance across these settings.
Figure 5: Visual results of Uni-NaVid in real-world. We deploy Uni-NaVid across diverse environments to execute instructions in a zero-shot setting. We provide third-person views with robot’s trajectory, showing effective navigation performance. We indicate the starting point as a blue dot and the ending point as a green arrow.
Method ScanQA
EM $\uparrow$ BLUE-1 $\uparrow$ ROUGE $\uparrow$ METEOR   $\uparrow$ CIDEr   $\uparrow$
V.N.+MCAN [105] 19.71 29.46 30.97 12.07 58.23
S.R.+MCAN [105] 20.56 27.85 30.68 11.97 57.56
3D-LLM(flamingo) [32] 23.2 32.60 34.80 13.5 65.6
NaviLLM [114] 26.27 39.73 40.23 16.56 80.77
BridgeQA [65] 31.29 34.49 43.26 16.51 83.75
Uni-NaVid 28.01 46.85 45.74 19.24 94.72
TABLE VIII: Embodied video question answering. Comparison on ScanQA [7] benchmark.
Method MSVD-QA MSRVTT-QA ActivityNet-QA
[…]
VideoLLaMA [107] 51.6 2.5 29.6 1.8 12.4 1.1
VideoChat [49] 56.3 2.8 45.0 2.5 26.5 2.2
VideoChatGPT [62] 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter [56] 67.5 3.7 57.0 3.2 45.7 3.2
Chat-UniVi [39] 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-VID [51] 69.7 3.7 57.7 3.2 47.4 3.3
VideoChat2 [50] 70.0 3.9 54.1 3.3 49.1 3.3
Video-LLaVA [53] 70.7 3.9 59.2 3.5 45.3 3.3
ST-LLM [57] 74.6 3.9 63.2 3.4 50.9 3.3
Uni-NaVid 69.6 3.9 59.3 3.5 51.4 3.7
TABLE IX: Video question answering. Comparison with leading methods (all based on Vicuna-7B [20]) on VQA benchamarks.
Figure 6: Vivusal results of Uni-NaVid on compostional tasks. The agent is required to execute complex instructions involving multiple navigation tasks. Our method successfully accomplishes these navigation tasks sequentially. Notably, both the instructions and environments are novel to our approach. Please refer to the supplementary videos.
Comparison on video question answering. We first evaluate our method on ScanQA [7] on Tab. VIII. Compared to mainstream baselines, we find that Uni-NaVid archives the best performance on four metrics, including BLUE-1 ($+17.9\%$), ROUGE ($+5.7\%$), METEOR ($+16.2\%$), and CIDEr ($+13.1\%$). This proves the superiority of our methods on spatial scene understanding. Note that the EM metric requires an exact match between the question and answer, which is not well-suited to our method, as it is designed to learn from diverse data and generate flexible responses.
We further evaluate our method on open-ended video question-answering benchmarks [13, 96, 10], as presented in Table IX. To ensure a fair comparison, we focus on methods that employ the same large language model backbone (Vicuna-7B [20]). The results indicate that even after extensive token merging (Sec. IV-B), Uni-NaVidachieves performance comparable to state-of-the-art methods. This demonstrates the effectiveness of both our token merging and training strategies, while also highlighting robust open-world understanding capabilities.
VI-C Qualitative Results in Real-World
We conducted extensive experiments on real-world environments (experiment details are provided in the supplemental material) under diverse environments in a zero-shot manner. Notably, both the instructions and environments are novel to our method. We first evaluated the performance of individual navigation tasks (Fig. 5), including (A) vision-and-language navigation, (B) object goal navigation, (C) embodied question answering, and (D) human following. We found that Uni-NaVid can understand diverse instructions and demonstrates impressive performance in long-horizon navigation tasks (e.g., navigating across hallways and entering rooms), as well as in searching for out-of-vision objects and answering subsequent questions. Moreover, the agent is capable of following a human even when the person’s appearance deviates from the description of the avatar in the human-following dataset (Sec. V-A). The statistics of the corresponding real-world experiments can be found in the Supplemental Material.
In addition to individual navigation tasks, we also evaluate our method on more complex instructions involving multiple navigation tasks (Fig. 6). In this scenario, the agent is required to sequentially execute the navigation tasks described in the language instructions. Our model demonstrates impressiove performance in aligning the current navigation process with the instructions to reason about the current state of navigation. Furthermore, we provide a detailed illustration of action prediction during navigation in Fig. 7, where we plot the predicted action probabilities of Uni-NaVid. Notably, with only slight differences in object descriptions, e.g., ’chair with a toy’ and ’chair with a sweater’. Specifically, our method successfully distinguishes between the locations and predicts actions accordingly. Interestingly, the action probabilities (for the next four actions) reveal a sequential order of actions: first turning right/left, followed by moving forward. We provide additional visual results of our method in the Supplemental Material and encourage the audience to view our video, which showcases the real-world performance of our method.
Figure 7: Action prediction on the VLN tasks. We evaluate Uni-NaVid on challenging open-vocabulary objects, requiring it to recognize the target objects and follow the specified motions. We provide the predicted action probabilities (for the next four actions) to demonstrate its break-in navigation capability.
Figure 8: Comparsion on multi-task training and data scale. (a) We present the multi-task synergy of our method, illustrating the performance comparison between training with a single task and training with multiple tasks; (b) we demonstrate the performance across different navigation tasks under varying numbers of training samples.
VI-D Ablation Study
Visualization of training strategy. We present a visualization of the training strategy’s performance in Figure 8. In Fig. 8 (a), we compare training on a single navigation task with training across multiple tasks. The results demonstrate the synergistic benefits of multi-task learning, which yields consistent performance improvements across all navigation tasks. Notably, VLN, ObjectNav, and EQA exhibit more significant improvements, while Following shows relatively smaller gains. We attribute this difference to the lower reliance of the Following task on historical context. Additionally, we investigate the influence of data scale on navigation performance (Figure 8 (b)). We observe that performance improves across all navigation tasks with larger data volumes. However, the incremental gain diminishes (from 3M to 6M samples), potentially due to limitations in the data diversity of simulators. Specifically, for the Following task, the reason for the slower convergence is the heavy occlusion caused by obstacles or other humans. This highlights the need for more high-quality following data samples, which can enable our model to learn more effectively and perform better in highly dynamic environments.
Ablation on training strategy and architecture. We conduct experiments to evaluate the effectiveness of the training strategy and token merging designs (Tab. X). Our results indicate that the absence of <NAV> and VQA data leads to a performance decline across all tasks, similar findings can be found in [14, 111]. Notably, the performance drop is most obviously in EQA, as the lack of <NAV> special token makes the model misinterpret whether it should answer questions or output actions. Additionally, without VQA data, the agent’s ability to answer questions drops significantly, almost rendering it incapable of correctly answering questions. We believe this is due to the catastrophic forgetting problem in LLMs, where the model loses open-world knowledge by being trained solely on navigation-related data.
From the performance of different memory designs, we find that both short-term and long-term memory visual tokens contribute to performance improvements. In particular, the VLN task shows the most significant performance drop ($-80.3\%$ SR) when visual memory is removed, as the lack of memory hinders the alignment of visual history with instructions. For the Following task, the absence of memory results in only a minor performance decline ($-8\%$ SR), as this task primarily relies on recent frames to track the target. Additional ablation studies on architecture and hyperparameters are provided in the Supplementary Material.
[…]
No $<$Nav$>$ token 35.2 69.1 20.4 55.1
No VQA data 40.5 50.6 1.19 58.8
Curr. 9.61 44.3 32.5 56.3
Curr.+Short. 39.7 67.8 44.1 59.7
Curr.+Short.+Long. 48.7 73.7 47.3 61.2
TABLE X: Ablation study on training strategy and architecture. For each ablation type, we retrain the entire model and evaluate its performance across four navigation tasks.
VII Limitations
Despite the promising results, Uni-NaVid has several limitations. First, Uni-NaVid is trained and evaluated on four well-defined navigation tasks, while there exists a large body of literature on insightful and practical navigation datasets [86, 112]. We believe that collecting data from these datasets could further enhance the navigation capabilities of our method. Second, our method is designed to acquire multi-task navigation capabilities under the assumption that the robot is of standard size (see Section V-A). To extend it to robots of general sizes, a convincing approach is to incorporate prior knowledge of the robot’s size, as demonstrated in [24, 29]. Third, our method is currently limited to predicting simple trajectories composed of a short horizon of future low-level discrete actions. This limitation could be alleviated by extending the moel to predict continuous and smooth trajectories with techniques from motion planning [79, 81] or autonomous driving [18, 52].
VIII Discussion and Conclusion
In this paper, we introduce an efficient vision-language-action (VLA) model, Uni-NaVid, designed to acquire general embodied navigation skills through learning multi-task navigation data. To efficiently encode the online-captured video sequences during navigation, we develop an online visual token merging mechanism that separately processes current observations, short-term observations, and long-term observations. This design enables our approach to operate at an average speed of 5 Hz. We also collect 3.6 million navigation data points across four highly demanded embodied navigation tasks, including vision-and-language navigation, object goal navigation, embodied question answering, and human following. Extensive experiments and ablation studies demonstrate that our method achieves SOTA-level performance using only monocular videos as input, highlighting our model’s superior capability in learning multiple navigation tasks. Moreover, we deploy Uni-NaVid in real-world environments, demonstrating impressive generalizability and versatile navigation performance in real worlds.
Future works. Our work serves merely as a starting point of general-purpose navigation, and we hope it will inspire future directions in this field:
  • Benchmarking. With the consistent development of embodied navigation, there is a growing need for general-purpose navigation benchmarking. Such a benchmark would help researchers better position their work and drive progress in the navigation community.
  • Architecture. We would like to further enhance the practicality of our architecture by tackling very long-horizon tasks (e.g., navigating across buildings) and incorporating advanced motion planning techniques [81, 18].
  • Application. We would like to apply our method to applications such as robotic guide dogs and home service robots. Additionally, we are excited to extend this technique to other embodied AI tasks, such as mobile manipulation [110, 92, 55].